• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jglister

    Log entry time 00:01:00 on August 31, 2006

    Entry number 184026

    keyword=Swing shift summary

    J.Glister (SL), A. Shahinyan (TO), M. Potokar (3rd)

    16:00 Made access to move 3/4" carbon doors out, also
    Mark Stevens reversing polarity of left arm magnets
    and setting momentum to 496.1 MeV/c
    17:00 SO in and turning fpp high voltages on
    17:23 Left arm at 103.4 degrees
    17:34 Back in beam permit, asked for 20uA, raster on
    17:55 Moved to 15cm LD2 target
    18:03 BPM jumping around, asked MCC to turn fast feedback on
    18:27 Moved to 15cm LH2 target
    18:45 Adjusted left arm D1 field reading to give proper momentum
    (496.1 MeV/c), re-doing rate check runs with new momentum
    19:18 Moved to 15cm LD2 target
    19:25 MCC reports our fast feedback is tripping their DAQ, shut off for
    now
    19:34 Starting with Gamma D production data: LD2, 5% radiator in
    21:06 MCC called to say they were taking beam away for 10-15 minutes
    for PZT tests
    21:21 MCC called: tests done & beam back
    22:25 Right arm coda died
    23:30 Right arm rebooted (finally)
    23:45 Right arm coda freezing again


    Runs

    Run# Target Radiator Type #Events/comment
    --------------------------------------------------
    3619 C out Cosmics 100k
    3620 C out Pointing 100k

    3621 LD2 out Rate Check 1M
    3622 LD2 in " 200k

    3623 LH2 in " junk
    3624 LH2 in " 200k
    3625 LH2 in " 200k (p corrected)
    3626 LH2 out " 200k

    3627 LD2 out " 200k
    3628 LD2 in " 200k
    3629 LD2 in GD Prod. 1.9M
    3630 " " " "
    3631 " " " "
    3632 " " " 1.3M (stopped early for PZT
    tests)
    3633 " " " 1.9M
    3634 " " " "
    3635 " " " "
    3636 " " " "
    3637 " " " "
    3638 " " " junk
    handing off 3639 to owl shift